OpenStudy (wasiqss):

5 - Homework - An example. This homework required some thought and work. If you manage to hand in some solution, this will improve your homework statistics. A student has a current income y1 and expects future income y2. He/She plans current consumption c1 and future consumption c2 in order to maximize the utility function U = ln(c1) + 1 1 +  ln(c2) c1, c2 > 0 where  is his/her discount rate. Borrowing or lending costs/pay an interest rate of r. Write down the consumption of the student in the future depending on the consumption today. This means try to think about a way to express

OpenStudy (wasiqss):

This means try to think about a way to express c2 in terms of c1 (Thats the first step) Now find the optimal borrowing plan (plug in, what you found above and maximize the utility with respect to c1).

OpenStudy (wasiqss):

Plz help this question got hell out of me!

OpenStudy (anonymous):

Can you clear up the notation on that utility function?

OpenStudy (wasiqss):

\[u=\ln(c1)+(\ln(c2)/(1+\delta)\]

OpenStudy (wasiqss):

c1,c2>0

OpenStudy (anonymous):

\[ U=\ln c_1+\frac{\ln c_2}{1+\delta} \]And \(\delta\) is the discount rate?

OpenStudy (wasiqss):

yes

OpenStudy (anonymous):

Out of curiosity, what course is this for?

OpenStudy (wasiqss):

BBA

OpenStudy (anonymous):

BBA?

OpenStudy (wasiqss):

Bachelors in Business administration

OpenStudy (wasiqss):

i want this one today anyhow :(

OpenStudy (wasiqss):

n i have no freaking clue how to attempt this :(

OpenStudy (anonymous):

That's funny, I have a Bachelor in Business Administration :P

OpenStudy (wasiqss):

lol, cool

OpenStudy (anonymous):

I mean, through pure manipulation you can do the following:\[ U=\ln c_1+\frac{\ln c_2}{1+\delta}\\ U-\ln c_1=\frac{\ln c_2}{1+\delta}\\ (U-\ln c_1)(1+\delta)=\ln c_2\\ e^{(U-\ln c_1)(1+\delta)}=c_2 \] That's \(c_2\) in terms of \(c_1\) and U.

OpenStudy (wasiqss):

now whats the next step

OpenStudy (anonymous):

I'm not sure, I don't think that's what they're actually looking for when they say \(c_2\) in terms of \(c_1\). The goal here is to have something where we can differentiate U wrt \(c_1\) and set it equal to zero to maximize the function.

OpenStudy (wasiqss):

now we have to differentiate wrt to c1

OpenStudy (wasiqss):

@nbouscal can you help in differentiating it as it is looking very complex

OpenStudy (wasiqss):

@across help me differentiate this wrt to c1 plz

OpenStudy (wasiqss):

@ash2326 @TuringTest plz help me differentiate it wrt to c1 plz

OpenStudy (wasiqss):

so ttired to think